OpenStudy (anonymous):

A certain medicine is given in an amount proportional to a patient's body weight. Suppose a patient weighing 100 pounds requires 125 milligrams of medicine. What is the amount of medicine required by a patient weighing 216 pounds?

OpenStudy (anonymous):

The patient is 216/100 times heavier than the 100 pounds patient so need 216/100 times the amount. Hopefully that's enough of a clue.

OpenStudy (anonymous):

Create a proportion: 100lb/125mg = 216lb/ x mg. Cross multiply; 100x=125*216 x= 270 mg

OpenStudy (anonymous):

125/100 x 216 = 270... i would say 270 milligraams of medicine:)

OpenStudy (anonymous):

Jay123 your right thats the answer THANKU

OpenStudy (anonymous):

No prob! I would say openstudy624 is right too:)

OpenStudy (anonymous):

whats ur email u know for future help

OpenStudy (anonymous):

ahem... i m not sure if i will be able to help u in future cuz that depends on the question... do u really need my email?

OpenStudy (anonymous):

most of what i do is algebra